OpenStudy (anonymous):

find the derivative of y=x^4+3sqrt(x)/x^2

OpenStudy (anonymous):

Use power rule

OpenStudy (anonymous):

\[ y=x^4+\frac{3\sqrt x}{x^2} = x^4+3x^{-3/2} \]

OpenStudy (anonymous):

uhm wait no thats not the problem

OpenStudy (anonymous):

it's x^4+3sqrt(x) all divided by x^2

OpenStudy (anonymous):

\[ y = \frac{x^4+3\sqrt x}{x^2}= \frac{x^4}{x^2} + \frac{3\sqrt x}{x^2} = x^2+3x^{-3/2} \]

OpenStudy (anonymous):

\[ dy = d(x^2+3x^{-3/2}) = d(x^2)+3~d(x^{-3/2}) \]

OpenStudy (triciaal):